OpenStudy (cookiemonster18):

1. you and a friend balls for like a salad and a small drink. between the two of you, you have $8.00. A salad costs $2.49 and a small drink is $0.99. Can either of you have a second salad or a drink? assume there is no sales tax. A.yes,one Salad B.yes,one of each C.yes,one drink D.No,you cannot 2.A student spends more than 2 hours on his math and English homework. If math takes about twice as long as enligh, what is tha maximum time that the student can spend on English? A. 1/3 hour B. 1/2 hour C. 1 hour D. 2/3 hour

Directrix (directrix):

I think that problem 2 above should begin as: A student spends NO more than 2 hours on his math. @CookieMonster18

Directrix (directrix):

What is the meaning of "balls" in problem one: you and a friend*** balls**** for like a salad

OpenStudy (cookiemonster18):

I think 1 is B and 2 is A am I correct? @Directrix

OpenStudy (texaschic101):

1. 8 - (2.49 + 0.99) = 8 - 3.48 = 4.52.....well, you and your friend can afford 1 of each 2. m + e > 2 m = 2e now we sub 2e + e > 2 3e > 2 e > 2/3.....the only one greater then 2/3 is C .

Directrix (directrix):

2 salads and 2 drinks cost 2*2.49 + 2*.99 = $ 6.96 You get 1 salad and 1 drink and the friend gets 1 salad and 1 drink. Of the $8.00, $6.96 has been spent. That leaves 8.00 - 6.96 = $1.04. You or the friend, but not both of you, could get 1 additional drink with $1.04,

Directrix (directrix):

A student spends NO more than 2 hours on his math and English homework. If math takes about twice as long as enligh, what is tha maximum time that the student can spend on English? Borrowing work from above, m + e 2 where m is math and e is english m = 2e now we sub 2e + e ≤ 2 3e ≤ 2 e ≤ 2/3 hrs. @CookieMonster18

Directrix (directrix):

>>I think 1 is B and 2 is A am I correct? I do not agree with either answer. @CookieMonster18

OpenStudy (cookiemonster18):

Well what would be the answer for 1 and 2 then? :/ because im having alot of trouble understanding because so many people are giving me different answers @Directrix I thought that maybe 1 was B and 2 was C but im unsure now. im very confused

Directrix (directrix):

I don't see people giving different answers. I see people providing work to help you find answers. I also see that both questions contain typos or omissions which caused confusion to the helpers.

Directrix (directrix):

On the first question, I posted this: 2 salads and 2 drinks cost 2*2.49 + 2*.99 = $ 6.96 You get 1 salad and 1 drink and the friend gets 1 salad and 1 drink. Of the $8.00, $6.96 has been spent. That leaves 8.00 - 6.96 = $1.04. *** You or the friend, but not both of you, could get 1 additional drink with $1.04.***** The starred sentence above makes the answer to #1 clear to me. Do you not see it? @CookieMonster18

Directrix (directrix):

On #2, the final line I wrote is this: e ≤ 2/3 hrs. If English study time is less than OR equal to 2/3 of an answer, what is the maximum study time for English?

Directrix (directrix):

How did you get option B for #1. >>B.yes,one of each Why C for #2? >>C. 1 hour

OpenStudy (cookiemonster18):

I got option B for #1 because a salad is $2.49 and a drink is $0.99 so if they both got 1 salad and 1 drink each it'll be (for the salad for both of them, $4.98) (for a drink for both of them i'll be, $1.98) total ($6.96) if they wanted another salad and drink each it would be another $6.96 but they only have $1.04 so that means they cant get any thing else because they BOTH want another drink but that would be ($1.98) but they only have $1.04 and they cant get another salad because that would be ($4.98) so the answer would be (D.No,you cannot) @Directrix

Directrix (directrix):

@CookieMonster18 I agree with what you posted. But, the question asks this: Can EITHER of you have a second salad OR a drink? With $1.04 left over, ONE person can have ONE more (a second) drink. That is why I selected option C for #1. C.yes,one drink In #2, e, the number of English study hours is e ≤ 2/3 hrs. The maximum value of e is less than or equal to 2/3 is 2/3. That is why I selected option D. 2/3 hour.